SPECTROSCOPIC SIGNATURES OF LARGE AMPLITUDE VIBRATIONAL MOTIONS IN H_{5}O_{2}^{+} AND H_{3}O_{2}^{−}

\underline{ANNE B. McCOY}, Department of Chemistry, The Ohio State University, Columbus, OH 43210; XINCHUAN HUANG, JOEL M. BOWMAN, Department of Chemistry and Cherry L. Emerson Center for Scientific Computation, Emory University, Atlanta, GA 30322; STUART CARTER, Department of Chemistry, University of Reading, Reading RG6 2AD, United Kingdom.

Recent spectroscopic studies of H_{5}O_{2}^{+}^{a}^{,}^{b}^{,}^{c} and H_{3}O_{2}^{−}^{d}^{,}^{e} in the 600 - 2000 cm^{−1} region have displayed strong transitions at frequencies that do not correlate to any of the expected harmonic vibrational frequencies in these systems. In this talk we will present results of calculations of the transition frequencies and intenisties using Diffusion Monte Carlo and variational calculations, performed using MULTIMODE.^{f} For these studies, we employ fully ab initio potential surfaces, recently reported by Huang, Braams and Bowman.^{g} Excellent agreement between the experimental and calculated spectra are obtained for both species. We find that the fact that the transitions frequencies are shifted by as much as several hundred cm^{−1} from the corresponding harmonic values is evidence of strong coupling among the zero-order normal mode vibrations as well as very large amplitude vibrational motions, even at relatively low levels of vibrational excitation.

aK. R. Asmis, N. L. Pivonka, G. Santambrogio, M. Brummer, C. Kaposta, D. M. Neumark, and L. Wöste, Science \underline{299}, 1375 (2003).

bT. D. Fridgen, T. B. McMahon. L. MacAleese, J. Lemaire, and P. Maitre, J. Phys. Chem. A \underline{108}, 9008 (2004).

cJ. M. Headrick, J. R. Roscioli, J. C. Bopp, and M. A. Johnson, J. Chem. Phys. \underline{121}, 11523 (2004)

dE. G. Diken, J. M. Headrick, J. R. Roscioli, J. C. Bopp, M. A. Johnson, A. B. McCoy, X. Huang, S. Carter, and J. M. Bowman, J. Phys. Chem. A \underline{109}, 171 (2005).

eE. G. Dinken, J. M. Headrick, J. R. Roscioli, J. C. Bopp, M. A. Johnson and A. B. McCoy, J. Phys. Chem. A \underline{109}, 1487 (2005).

fJ. M. Bowman, S. Carter, and X. Huang, Int. Rev. Phys. Chem. \underline{22}, 533 (2003).

gX. Huang, B. Braams, and J. M. Bowman, J. Chem. Phys. \underline{122}, 044308 (2005).